1. Kaya is riding her dirt bike eastward on a dirt road. She spots a pothole ahead. Kaya slows her bike from 14.0 m/s to 5.5 m/s in 6.0 s. What is her acceleration? 
1
Expert's answer
2022-03-16T18:28:23-0400

The acceleration

"a=\frac{\Delta v}{\Delta t}=\frac{5.5-14}{6.0}=-1.4\:\rm m/s^2"